OpenStudy (anonymous):

CAN SOMEONE HELP ME OUT PLEASE !!!!! solve for x: 12x – 17 = -41

OpenStudy (mertsj):

Add 17 to both sides first.

OpenStudy (anonymous):

12x - 17 = -41 Add 17 to both sides 12x = -24 Divide both sides by 12 x = -2

OpenStudy (anonymous):

so how would it look step by step

OpenStudy (anonymous):

12x - 17 = -41 +17 +17 12x = -24 /12 /12 x = -2

OpenStudy (anonymous):

Can you guys help me out with this one 4( x – 2 ) = 5( x + 3) and this one ? 5(x – 8) = -10

OpenStudy (anonymous):

Distribute the 4 and 5 giving you 4x - 8 = 5x + 14 +8 +8 4x = 5x +22 -5x -5x -x = 22 Divide by a -1 to both sides to get a positive x answer x = -22

OpenStudy (anonymous):

Second one- Again distribute the 5 giving you 5x - 40 = -10 +40 +40 5x = 30 /5 /5 x = 6

OpenStudy (anonymous):

thank you for helping me out

OpenStudy (anonymous):

No Problem!